OpenStudy (anonymous):

Simplify 4k2(3k3 + 5k2 + 6k – 7)

OpenStudy (callisto):

I'm not sure.. I think it's expand the expression. \[4k^2(3k^3 + 5k^2 + 6k – 7) \]\[=4k^2(3k^3) + 4k^2(5k^2) + 4k^2(6k) – 4k^2(7) \]\[=12k^{2+3} + 20k^{2+2} + 24k^{2+1} – 28k^2 \]\[=...?\]

OpenStudy (anonymous):

these are the choices hun :) 12k3 + 20k2 + 24k – 28 12k5 + 20k4 + 24k3 – 28k2 7k3 + 9k2 + 10k – 3 7k5 + 9k4 + 10k3 – 3k2

OpenStudy (callisto):

*it's about expanding

OpenStudy (callisto):

Perhaps you can get the answer after you do the addition for the power!~

OpenStudy (anonymous):

OMG its b thank you muah i love you i love you i love you i love you

OpenStudy (callisto):

:O Welcome.....